OpenStudy (anonymous):

Find an explicit formula for the sequence defined recursively by \[a _{1}=1\] and \[a _{k}=2a _{k-1}+6\] Then prove the formula is correct using mathematical induction

OpenStudy (anonymous):

First 5 terms (a_1,a_2,a_3,a_4,a_5) are 1,8,22,50,106

OpenStudy (across):

That's right.

OpenStudy (anonymous):

\[a _{2}=2a _{1}+6\rightarrow2(1)+6\rightarrow2^{1}+6\]

OpenStudy (across):

Can you come up with a formula for it?

OpenStudy (anonymous):

\[a _{3}=2a _{2}+6\rightarrow2(2+6)+6\rightarrow2^{2}+2*6+6\]

OpenStudy (anonymous):

Cant seem to, no

OpenStudy (anonymous):

my book says to write them in exponential when you can, which is why i am doing 2^2 instead of 2*2

OpenStudy (anonymous):

so I have that form for all the way up to a_5

OpenStudy (anonymous):

also, the difference between each term is 1,14,28,56

OpenStudy (anonymous):

so i am thinking maybe something to do with 7?

OpenStudy (anonymous):

\[a _{5}=2a _{4}+6=2(2^{3}+2^{2}*6+2*6+2*6+6)+6=2^{4}+2^{3}*6+2^{2}*6+2*6+6\]

OpenStudy (anonymous):

\[a _{5}=2a _{4}+6=2(2^{3}+2^{2}*6+2*6+2*6+6)+6=2^{4}+2^{3}*6+2^{2}*6+2*6+6\]

OpenStudy (anonymous):

Seeing any patterns?

OpenStudy (across):

You're doing a brilliant job. Do you notice the pattern with the 2, the 6 and the 7?

OpenStudy (anonymous):

We have 2^(n-1) in front everytime...

OpenStudy (across):

There's a formula to determine these series when you find out these patters. I forgot what it was called, but assuming that we're starting our count at 1, you can express it as something like\[\frac{a\cdot2^n+b}{c}.\]

OpenStudy (anonymous):

Sum of geomretic series?

OpenStudy (anonymous):

What is a,b,and c?

OpenStudy (anonymous):

Any tips mr.math?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

are there any exponents?

OpenStudy (mr.math):

I think so.

OpenStudy (anonymous):

I still dont see it for sure im pretty bad at detecting these

OpenStudy (anonymous):

obviously the first term is doubling somehow , 1-> 2 -> 4 ->8 -> 16

OpenStudy (mr.math):

I'm having a problem myself. I think there could be a better way to do it.

OpenStudy (mr.math):

Give me a minute!

OpenStudy (mr.math):

Here: \[a_1=1=1+(2^0-1)(7)\] \[a_2=1+7=1+(2^1-1)(7)\] \[a_2=1+3(7)=1+(2^2-1)(7)\] \[a_3=1+7(7)=1+(2^3-1)(7)\] ...

OpenStudy (anonymous):

(2^(n-1)-1)7

OpenStudy (mr.math):

I think you can figure it out now. If I find a better method to solving it, I'll let you know.

OpenStudy (mr.math):

Yeah, \(\large a(n)=1+7(2^{n-1}-1)\).

OpenStudy (anonymous):

Yea, i guess its all just manulation till its in a form where a patern is more recognizable

OpenStudy (anonymous):

Have to go now, ill be around later to make sense of this, thanks a lot

OpenStudy (mr.math):

It can be simplified into \[a(n)=7(2^{n-1})-6.\]

OpenStudy (mr.math):

You're welcome.

OpenStudy (anonymous):

How did you figure this one out?

OpenStudy (anonymous):

Was the solution found via just pure math magic manipulation, where you just have to see it. Or was some formula involved or something?

OpenStudy (zarkon):

there are techniques to solving these and formulas that one can use. This one is easy enough that you can build up a solution 1 2*1+6=2+6 \[2(2+6)+6=2^2+2\times 6+6=2^2+6(2+1)\] \[2(2^2+6(2+1))+6=2^3+2\times6(2+1)+6\]\[=2^3+6\times(2^2+2)+6=2^3+6\times(2^2+2+1)\] ... \[2^{n-1}+6(2^{n-2}+2^{n-2}+\cdots+2^2+2+1)\] \[=2^{n-1}+6\frac{2^{n-1}-1}{2-1}=2^{n-1}+6(2^{n-1}-1)=2^{n-1}+6\cdot2^{n-1}-6\] \[=2^{n-1}(1+6)-6=2^{n-1}(7)-6=7\cdot2^{n-1}-6\]

OpenStudy (anonymous):

Im trying to figure out how you got from 3rd line up to 2nd line up...

OpenStudy (zarkon):

geometric partial sum... \[\sum_{k=0}^{n-1}a^k=\frac{a^n-1}{a-1}\]

OpenStudy (anonymous):

hmm, not familiar

OpenStudy (anonymous):

anything to do with this theorem: http://imgur.com/V6Nnz

OpenStudy (zarkon):

no.... http://en.wikipedia.org/wiki/Geometric_series#Formula

OpenStudy (anonymous):

ok, ill read up on that

OpenStudy (anonymous):

But getting to do was a guess right

OpenStudy (anonymous):

based on the pattern

OpenStudy (anonymous):

Also, do you 2^n-2,2^n-2 on purpose or is it to be 2^n-2,2^n-1

OpenStudy (anonymous):

Ah ok, i found that theorem in my book: http://imgur.com/A1g9J

OpenStudy (zarkon):

that is a typo ... should be \[2^{n-1}+6(2^{n-2}+2^{n-3}+\cdots+2^2+2+1)\]

OpenStudy (anonymous):

should k not be 1 on the bottom of that summation?

OpenStudy (anonymous):

we go from 1

OpenStudy (zarkon):

k=0... \[2^0=1\]

OpenStudy (anonymous):

We are only given a_1 though

OpenStudy (anonymous):

I think I understand it all now, thats everyone, i <3 openstudy